Minutes RF structure meeting 09.08.2007

F. Gerigk, M. Pasini, S. Ramberger, M. Vretenar

news & facts:

  • high-power test stand in SM18 will be available at least until end of August,
  • CERN CCDTL prototype
    • now closed and under vacuum, vacuum seems to be ok,
    • the drift tube in cell 1 has been repaired (opened, cooling circuit repaired, closed with a new nose piece),
    • the copper from the tank covers has been removed in an acid bath and then they were copper plated including the area of the Helicoflex joints, the area around the joints on the tank side are still not covered with copper,
    • the frequency in all cells (in air) is now 352.335 MHz,

  • Russian CCDTL prototype
    • high-power tests should continue until we can measure the water flow and pressure,
    • according to Alexej the water flow in one drift tube should be around 17-22 liters to achieve the nominal cooling for the SPL duty cycle, so far we have probably 17-22 liters for the whole prototype,

new actions

Assigned to Start date Description State Result
MatteoPasini 2007-08-09 find frequency shift for CCDTL prototype, when going from air to vacuum and power for Linac4 and SPL duty cycle, compare tuned frequencies with Wanglers formulas CCDTL prototypes: test results, M. Pasini, M. Vretenar, R. Wegner, CARE-Report-2007-036-HIPPI edit
FrankGerigk 2007-08-09 ask Yannic for a pressure meter and compare water flow values with old estimates for SPL pressure meters are installed edit

updated ActionListRF

minutes by Frank Gerigk

-- FrankGerigk - 09 Aug 2007

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2008-06-30 - FrankGerigk
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright